# Samuel Hildebrand    Curtis Taylor    Sean Oesch    James M. Ghawaly Jr.    Amir Sadovnik    Ryan Shivers    Brandon Schreiber    Kevin Kurian

### v1

@@ II-A General-Purpose Multimodal Benchmarks
Early efforts to evaluate reasoning capabilities across diverse tasks include the MMLU-Pro benchmark by Wang et al., which provides a robust and challenging framework for assessing broad knowledge and reasoning in language models [1]. Expanding to multimodal settings, Yue et al. introduce MMMU, a comprehensive multimodal understanding suite aimed at expert-level evaluation across multiple academic disciplines [2]. Fan et al. propose NPHardEval4V, which dynamically generates NP-hard combinatorial puzzles paired with multimodal contexts to probe model adaptability [3].
@@ II-B Specialized Multimodal Benchmarks
Beyond general-purpose evaluations, several benchmarks target specific reasoning challenges. Kil et al. present MLLM-CompBench, measuring reasoning performance of multimodal LLMs on tasks from visual question answering to diagram interpretation [4]. Guo et al. introduce R-Bench, emphasizing graduate-level, cross-disciplinary problems spanning mathematics, biology, and engineering [5]. EMMA by Hao et al. evaluates reasoning over annotated image-text pairs, focusing on commonsense and factual inference in visual contexts [6].
@@ II-C RAG System Evaluation and Hallucination Detection
Yang et al. introduce the CRAG benchmark, which evaluates RAG systems’ generation capabilities under various retrieval conditions. However, CRAG focuses primarily on how generation behaves given retrieved context, rather than evaluating the retrieval mechanisms themselves or providing end-to-end pipeline assessment [7].
RAGAS (Retrieval Augmented Generation Assessment) provides a framework for evaluating individual RAG components through metrics such as context precision, context recall, and answer relevancy [8]. While useful for component-level analysis, RAGAS does allow for evaluation of RAG pipelines as a whole.
Detecting hallucinations in model outputs has been approached from several directions. Recent work includes probabilistic uncertainty estimation through semantic entropy (Farquhar et al.) and model-internal state analysis for real-time detection (Su et al.) [9, 10]. While effective, these methods often require access to token-level probabilities, hidden states, or repeated sampling—conditions that are impractical when evaluating retrieval pipelines as whole, or impossible in many cases when using closed-source RAG pipelines.
@@ III-A Dataset Construction
To construct our evaluation dataset, we curated 93 non-trivial questions and associated answers from a collection of 8 research papers in the fields of AI and machine learning.[11, 12, 13, 14, 15, 3, 16, 17] To answer these questions requires referencing various modalities within the 8 papers. The questions are organized into 5 categories:

### v2

@@ II-A General-Purpose Multimodal Benchmarks
Early efforts to evaluate reasoning capabilities across diverse tasks include the MMLU-Pro benchmark by Wang et al., which provides a robust and challenging framework for assessing broad knowledge and reasoning in language models [NEURIPS2024_ad236edc]. Expanding to multimodal settings, Yue et al. introduce MMMU, a comprehensive multimodal understanding suite aimed at expert-level evaluation across multiple academic disciplines [yue2024mmmu]. Fan et al. propose NPHardEval4V, which dynamically generates NP-hard combinatorial puzzles paired with multimodal contexts to probe model adaptability [fan2024nphardeval4vdynamicreasoningbenchmark].
@@ II-B Specialized Multimodal Benchmarks
Beyond general-purpose evaluations, several benchmarks target specific reasoning challenges. Kil et al. present MLLM-CompBench, measuring reasoning performance of multimodal LLMs on tasks from visual question answering to diagram interpretation [kil2024mllm]. Guo et al. introduce R-Bench, emphasizing graduate-level, cross-disciplinary problems spanning mathematics, biology, and engineering [guo2025rbenchgraduatelevelmultidisciplinarybenchmarks]. EMMA by Hao et al. evaluates reasoning over annotated image-text pairs, focusing on commonsense and factual inference in visual contexts [hao2025mllmsreasonmultimodalityemma].
@@ II-C RAG System Evaluation and Hallucination Detection
Yang et al. introduce the CRAG benchmark, which evaluates RAG systems’ generation capabilities under various retrieval conditions. However, CRAG focuses primarily on how generation behaves given retrieved context, rather than evaluating the retrieval mechanisms themselves or providing end-to-end pipeline assessment [NEURIPS2024_1435d2d0].
RAGAS (Retrieval Augmented Generation Assessment) provides a framework for evaluating individual RAG components through metrics such as context precision, context recall, and answer relevancy [es2023ragas]. While useful for component-level analysis, RAGAS does allow for evaluation of RAG pipelines as a whole.
Detecting hallucinations in model outputs has been approached from several directions. Recent work includes probabilistic uncertainty estimation through semantic entropy (Farquhar et al.) and model-internal state analysis for real-time detection (Su et al.) [farquhar2024detecting, su2024unsupervisedrealtimehallucinationdetection]. While effective, these methods often require access to token-level probabilities, hidden states, or repeated sampling—conditions that are impractical when evaluating retrieval pipelines as whole, or impossible in many cases when using closed-source RAG pipelines.
@@ III-A Dataset Construction
To construct our evaluation dataset, we curated 93 non-trivial questions and associated answers from a collection of 8 research papers in the fields of AI and machine learning.[wang2023docllmlayoutawaregenerativelanguage, utkarsh2025physicsconstrainedflowmatchingsampling, oesch2024pathautonomouscyberdefense, yang2025chartmimicevaluatinglmmscrossmodal, wang2025enigmaevalbenchmarklongmultimodal, fan2024nphardeval4vdynamicreasoningbenchmark, zhmoginov2022hypertransformermodelgenerationsupervised, zhang2024multimodalchainofthoughtreasoninglanguage] To answer these questions requires referencing various modalities within the 8 papers. The questions are organized into 5 categories: